OpenStudy (anonymous):

Please HELP! f(x)= 2x+3 for x less than equal to 0 g(x)= x^2-6x for x less than equal to 3 Find the set of values of x which satisfy gf(x) is less than equal to 16.

OpenStudy (anonymous):

I got x less than equal to -/+ 2.5. But the answer says the range is -2.5<x<0. How so?

OpenStudy (anonymous):

I don't know , give me your work, and we check together, is it ok to you?

OpenStudy (anonymous):

Okay. Sure.

OpenStudy (anonymous):

So, I did it like this. gf(x)= (2x+3)^2-6(2x+3) =4x^2-9

OpenStudy (anonymous):

Then I put \[4x^2-9\le16\]

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

\[x \le2.5\]

OpenStudy (anonymous):

So after putting on the number line, the range I got was

OpenStudy (anonymous):

+- 2.5

OpenStudy (anonymous):

\[-2.5\le x \le 2.5\]

OpenStudy (anonymous):

I know what the mistake is. you have to add the condition of f (x) which is x<0

OpenStudy (anonymous):

Why not consider g(x) x <3 ? That would allow x=2.5?

OpenStudy (anonymous):

because if x <0 , of course x<3 , right?

OpenStudy (anonymous):

satisfy that condition already

OpenStudy (anonymous):

you have 3 conditions: x<0 x<3 and -2.5< x < 2.5 combine them you have to take the answer that -2.5<x<0

OpenStudy (anonymous):

because for that combination, you satisfy 2 conditions above.

OpenStudy (anonymous):

got what i mean?

OpenStudy (anonymous):

Okay. Thank you so much! :)

OpenStudy (anonymous):

yw